OpenStudy (anonymous):

ok i need help if x^2 = y^2 + 39 and x = y +3 this is a substitution method please help

hartnn (hartnn):

What have you already tried ?

OpenStudy (anonymous):

(y + 3)^2 = y^2 = 39 Y^2 + 3y + 3y +3y + 9 = y^2 + 39 That is it so far

hartnn (hartnn):

check (y+3)^2 again \((a+b)^2 =a^2+2ab+b^2\)

hartnn (hartnn):

there will be only 2 "3y"s and not 3

OpenStudy (anonymous):

could you draw out the corrected step for me please

hartnn (hartnn):

(y+3)^2= y^2 + 39 y^2 +3y+3y+9 = y^2 +39 good so far ?

OpenStudy (anonymous):

k wait

OpenStudy (anonymous):

i have to move all of them to one side to equal 0

hartnn (hartnn):

you can...but our main aim is to find 'y' so you can move all y terms on one side and all constants(numbers) on other

OpenStudy (anonymous):

ok so s it ok if i move both 3ys and the 9 and then subtract the y^2 and be left with 0 = 6y + 30 is that ok or no?

hartnn (hartnn):

if you're moving both 3y's to other sign...means you're SUBTRACTING those 2 3y's from both sides, right ? so wouldn't it be 0=-6y+30 ?

OpenStudy (anonymous):

oh yeah and then what cause im trying to find 2 ys

OpenStudy (anonymous):

sry if im onfusing you

hartnn (hartnn):

so 6y =30 ok ?

OpenStudy (anonymous):

ok let me give you an example because i think im not asking the question right

OpenStudy (anonymous):

x^2 - 4x + 4 = 0 (x- 2) (x-2) x = 2, 2 that is what im trying to do

hartnn (hartnn):

see these steps, you may get it y^2 +3y+3y+9 = y^2 +39 y^2 -y^2 + 6y = 39-9 0 +6y = 30 y= 30/6 = .. any doubt in any step ?

OpenStudy (anonymous):

so if y= 5 would that be the only answer

hartnn (hartnn):

find the corresponding x value

hartnn (hartnn):

and yes, that x,y pair will be the only solution :)

OpenStudy (anonymous):

ok so one of my equations is x = y + 3 x = 5 + 3 x = 8

hartnn (hartnn):

yep, 8,5 is the only solution

OpenStudy (anonymous):

now i know the first is right but the thing is i must have two ys and xs how would i get the other ones?

OpenStudy (anonymous):

oh give me a moment

hartnn (hartnn):

its not necessary to have 2 pairs of solutions

OpenStudy (anonymous):

omg thank you so much

OpenStudy (anonymous):

i want to cry right now

hartnn (hartnn):

http://www.wolframalpha.com/input/?i=x%5E2+%3D+y%5E2+%2B+39+and+x+%3D+y+%2B3+ hyperbola : x^2 = y^2 + 39 and line : x = y +3 meet at only one point and hence there is only 1 solution don't cry please :P

hartnn (hartnn):

you're most welcome ^_^ since you're new here \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

is that program free?

hartnn (hartnn):

thats just a website link, it won't charge you anything for its basic functionality

OpenStudy (anonymous):

thank you i am now your fan i was gonna struggle on that one if you didn't show me

hartnn (hartnn):

happy to help :)